# A theory of ecological invasions and its implications for eco-evolutionary dynamics

### v1

@@ H.4 Eco-evolutionary dynamics of MiCRM
where $\rho$ is the correlation between mutant and parent traits, and $z_{c}$ is a random vector composed of uncorrelated standard normal variables. For our analysis, we used $\mu_{c}=10$ and $\rho=0.8$. After we sampled the invader and added it to the community, we solved for the new ecological steady-state using the iterative optimization algorithm common to all MiCRM simulations. We removed extinct species if they were below a threshold abundance of $0.1$ (chosen arbitrarily). We repeated this iterative eco-evolution until we observed 40 successful invasions in each community.
In figure 3(b), we applied hierarchical clustering to the consumer preference matrix to group consumers and resources based on their pairwise correlations. We computed Pearson correlation coefficients for rows (consumers) and columns (resources) separately, and performed clustering using the average linkage method. We reordered the matrix based on the resulting dendrograms to highlight correlation of similar consumers and resources. We visualized the reordered matrix as a heatmap with interaction strengths represented by color intensity.
In Figure 3(c), we calculated the average extinction probability versus consumer abundance by grouping the consumers into bins according to their pre-invasion abundance and computing the average extinction probability for consumers within each bin. The bins are grouped with approximately the same number of samples per bin. Number of bins in this plot is chosen to be 13, with around 350 samples per bin for the evolved systems, and around 460 for the assembled systems. The error bars represent the standard error for both axes, calculated for each bin.
Appendix I Methods for Data Analysis and Visualization
@@
We repeated the same set of procedures for these 3 different datasets as follows: (1) we first took the median abundance of each species in the largest ecosystem (containing all species) as the true values corresponding to new community abundance $X_{S}^{\text{new}}$ and invader abundance $X_{0}$, plotted on the y-axis; (2) then we considered the abundance of smaller ecosystems with one fewer species than the largest system and treated the missing species as an invader to those smaller systems as $X_{S}^{\text{old}}$; (3) Using our framework described by Eq. (2)-(7), we predicted $X_{S}^{\text{new}}$ and $X_{0}$. Since for each dataset, researchers repeated the same experiment with the same community composition 5 to 10 times, we had a prediction for each of these replicas. In Fig. 5, we plotted the median values on x-axis, and used the 25th and 75th percentiles as error bars.

### v2

@@ Appendix I Algorithm to solve the self-consistency equations
where $\rho$ is the correlation between mutant and parent traits, and $z_{c}$ is a random vector composed of uncorrelated Gaussian variables with zero mean and standard deviation $\sigma_{c}/\sqrt{n}$. For our analysis, we used $\mu_{c}=10$ and $\rho=0.8$. After we sampled the invader and added it to the community, we solved for the new ecological steady-state using the iterative optimization algorithm common to all MiCRM simulations. We removed extinct species if they were below a threshold abundance of $0.1$. We repeated this iterative eco-evolution until we observed 40 successful invasions in each community.
In Fig. 4(b), we applied hierarchical clustering to the consumer preference matrix to group consumers and resources based on their pairwise correlations. We computed Pearson correlation coefficients for rows (consumers) and columns (resources) separately, and performed clustering using the average linkage method. We reordered the matrix based on the resulting dendrograms to highlight correlation of similar consumers and resources. We visualized the reordered matrix as a heatmap with interaction strengths represented by color intensity.
In Fig.4(c), we calculated the average extinction probability versus consumer abundance by grouping the consumers into bins according to their pre-invasion abundance and computing the average extinction probability for consumers within each bin. The bins are grouped with approximately the same number of samples per bin. Number of bins in this plot is chosen to be 13, with around 350 samples per bin for the evolved systems, and around 460 for the assembled systems. The error bars represent the standard error for both axes, calculated for each bin.
K.5 Consumer resource model with (Monod) Type-II functional response
In Fig. 2 (i)–(k), and Fig. S3(e)–(h), we simulated an ensemble of 640 consumer-resource models employing Monod (Type-II) functional responses. The pre-invasion communities were sampled from a regional species pool consisting of 30 consumers, with the number of resources set to $n=30$. The Monod constant $h$ was chosen as $h=3$, comparable in magnitude to the resource abundances, thereby introducing a moderate degree of nonlinearity. Consumer preferences were sampled similarly to other models, with parameters $\mu_{c}=3$ and $\sigma_{c}=1$. Environmental variables were held constant across species, with values $k=3$ and $u=1$.
In Fig. S5, we varied $h$ to sample models with different non-linearity. We choose $10$ values of $h$ linearly from $h=3$ to $h=10$. For each value we sampled $100$ systems. Other parameters are kept the same as before. All simulations for this model is performed with numerical ODE simulator ’LSODA’, with parameters atol=1e-11, rtol=1e-11, and we checked convergence by ensuring the maximum derivative was less than 1e-10.
Appendix L Methods for Data Analysis and Visualization
@@
We repeated the same set of procedures for these 3 different datasets as follows: (1) we first took the median abundance of each species in the largest ecosystem (containing all species) as the true values corresponding to new community abundance $X_{S}^{\text{new}}$ and invader abundance $X_{0}$, plotted on the y-axis; (2) then we considered the abundance of smaller ecosystems (assemblages) with exactly one fewer species than the largest system and treated the missing species as an invader to those smaller systems as $X_{S}^{\text{old}}$; (3) Using our framework described by Eq. (2)-(7), we predicted $X_{S}^{\text{new}}$ and $X_{0}$. Since for each dataset, researchers repeated the same experiment with the same community composition (assemblages) 5 to 10 times, we had a prediction for each of these replicas. In Fig. 5, we plotted the median values on x-axis, and used the 25th and 75th percentiles as error bars.
Appendix M Additional experimental data
To further assess the applicability of our framework to experimental systems, we present results for three additional sets of biodiversity-ecosystem functioning (BEF) experimental data [44, 45, 46]. We used Lotka-Volterra interaction matrices fitted in prevoiusly work [43], where phylogenetic relationships are assumed to shape species competition. Among the four models fitted with similar performance, we selected Model 3, which provides an intermediate level of complexity. Each dataset comprised species assemblages from distinct ecosystems. Compared to the datasets analyzed in the main text, we found two key differences. First, most assemblages lacked replicates, preventing us from estimating variability or error bars. Second, the datasets did not include exhaustive combinations of all possible species; consequently, pairs of assemblages used for comparison may differ by multiple species, with several considered as invaders or extinctions.
To align with the assumptions of our framework, we filtered pairs of assemblages based on “similarity”, defined as the fraction of shared species relative to the total in either assemblage. In Fig. S7, to enable comparable analyses to the main text, we first limited to systems with a clearly defined single invader, then applied different similarity thresholds to show a comparable number of invasion results on the plots. For (a), covering 9 years of the Biodiversity II experiment [44] and 9 years of the Wageningen experiment [45], a similarity threshold of 0.85 was applied. For (c), corresponding to the Cadotte 2013 dataset [46], we used a threshold of 0.6. Additionally, in Fig. S8, we plotted all pairs of assemblages with a similarity above 0.5, coloring points by similarity value. There is a clear trend showing that the similarity value positively correlates with the prediction performance, meaning that pairs of assemblages with higher species overlap tend to yield more accurate predictions within our framework.
Figure S7: Theory applicable to bio-diversity ecosystem and functioning (BEF) experiments. Scatter plots comparing experimentally measured abundances with theoretical predictions for three datasets of assemblages of plant ecosystems. We considered pairs of assemblages which differed from each other by the introduction of exactly one species which we call the invader, and variable numbers of extinctions events. Using our theory for the GLV model on these datasets, we predict both the invader abundance (left) and post-invasion community abundances (right). There is no replicates in general and therefore no error bars. Both axes use a sym-log scale, with linear scaling applied below the smallest non-zero data point. The datasets are all plant communities from different experiments, filtered to pairs of assemblages with high species similarity (defined as ratio of shared species in two assemblages) for visual clarity, (a) 9 years of Biodiversity II experiment [44] with similarity larger than 0.85, (b) 9 years of Wageningen experiment[45] with similarity larger than 0.85 and (c) one dataset Cadotte 2013 with similarity larger than 0.6. See Fig. S8 for all pairs of assemblages with similarity larger than 0.5 [46].
Figure S8: Extended version of Figure S7 with variable number of invaders and number of extinction events. The marker color indicates the similarity of species between the pairs of assemblages, defined as the number of shared species divided by the total number of species in these assemblages. Only assemblages with more than half of shared (similarity>0.5>0.5) species are included. The plots are (a) 9 years of Biodiversity II experiment [44] with (b) 9 years of Wageningen experiment [45], and (c) one dataset Cadotte 2013 [46].
Figure S9: Dressed invasion fitness correlates more strongly with invader abundance in experimental data. Scatter plots comparing experimentally measured invader abundances with invasion fitness and dressed invasion fitness, which are both computed from the fitted LV models and the pre-invasion abundance. Points indicate median value for measurements, with error bars indicating variability across replicates. Both axes use a sym-log scale, with linear scaling applied below the smallest non-zero data point. The datasets span (a)–(b) plant communities [39], (c)–(d) herbivore–algae communities [40], and (e)–(f) ciliated protist communities [41].
Figure S10: Additional data of dressed invasion fitness correlates more strongly with invader abundance in experimental data. Comparison of the relationship between invader abundance and both dressed invasion fitness and invasion fitness, shown in analogous plots to Fig. S9, for additional plant experiments: (a)–(b) nine years of the Biodiversity II experiment [44] with similarity larger than 0.85, (c)–(d) nine years of the Wageningen experiment [45] with similarity larger than 0.85, and (e)–(f) one dataset from Cadotte 2013 with similarity larger than 0.6. See Fig. S8 for all pairs of assemblages with similarity larger than 0.5 [46].
Figure S11: Extinction probability increases with invader-resident correlation, consistent with competition exclusion. We used the same simulation results from the MiCRM model in Fig. 4 to plot how the extinction probability of a resident species depends on its phenotypic similarity with an invader. Similarity is measured using correlation between consumption preferences of residents and invaders. For evolved communities, the highest correlation would be between parents (residents) and their mutants (invaders). The plot shows that resident species are indeed more likely to become extinct when their consumer preferences are more correlated with an invader, consistent with competitive exclusion [48, 43]. However, note that even at high correlation, extinction probability is still $\sim 60\%$. This indicates that in roughly 40% of cases, invaders can still coexist with highly similar species as long as they are embedded in a diverse community.